OpenStudy (anonymous):

The faces of an 8-sided die are numbered from 1 to 8. What is the probability of rolling a 5 three times in ten rolls?

OpenStudy (kropot72):

\[P(three\ 5's\ out\ of\ 10\ rolls)=10C3\times (\frac{1}{8})^{3}\times (\frac{7}{8})^{7}\]

OpenStudy (anonymous):

Okay! You dont have to do it for me, but how do I go about doing the multiplication for 1/8 to the power of three?

OpenStudy (anonymous):

Would it look like this? P(x=3) = \[10C3(0.1)^{3}(0.7)^{8} \] ?

OpenStudy (kropot72):

Turn the fractions into decimals: 1/8 = 0.125 7/8 = 0.875 So the required probability is given by the following: \[120\times (0.125)^{3}(0.875)^{7}\]